OpenStudy (anonymous):

If f(x) = 2x − 4 and g(x) = 1 over 2x − 2, find f[g(−4)]. 12 −12 −8 8

OpenStudy (anonymous):

I'm not sure I understand

OpenStudy (anonymous):

neither do i.

OpenStudy (anonymous):

I can't really tell what you are trying to ask, maybe you could write it out again?

OpenStudy (anonymous):

\[g(x)=\frac{ 1 }{ 2x-2 }\]Is that what you mean?

OpenStudy (anonymous):

Well if that is the case then\[g(-4)=\frac{ 1 }{ 2(-4)-2 }=\frac{ -1 }{ 10 }\]

OpenStudy (anonymous):

Now all that is left is to put that number in to the f(x) and then job done!

OpenStudy (anonymous):

 this is exactly how it looks in the lesson.

OpenStudy (anonymous):

Are there any brackets around anything in the g(x)?

OpenStudy (anonymous):

nope. there is just the parentheses around the x

OpenStudy (anonymous):

@hea

OpenStudy (anonymous):